what is the domain of y=sq.root of (9-x^2) i htink the answer may be x is less than or equal to 3 that is what i got when i worked it out
Found 2 solutions by stanbon, AnlytcPhil:
Answer by stanbon(48535)

what is the domain of y=sq.root of (9-x^2)
-----------------------
9-x^2 must be greater than or equal to zero.
9-x^2>=0
(3-x)(3+x)>=0
Put 3 and -3 on a number line and figure out where
the rest of the solution set lies.
There are three intervals on the line. Check each interval, as follows:
Interval I is (-infinity,-3), so check with x=-10.
You get (3-(-10))(3+(-10)) which is <0 so no solutions there.
Interval II is (-3,3), so check x=0.
(3-0)(3+0) is >0 so this interval is part of the solution set.
Interval III is (3,+infinity), so check with x=10.
You get (3-10)(3+10) which is <0 so no solutions there.
SOLUTION:
All x values between -3 and +3 including 3 and -3: or -3<=x<=0
Cheers,
Stan H.

Answer by AnlytcPhil(1116)


What is the domain of y=sq.root of (9-x^2) 
I think the answer may be x is less than 
or equal to 3 that is what I got when I 
worked it out

No that is incorrect.

For
            ____
       y = √9-x²

we require that what's under the radical
not be negative, that is, greater than
or equal to 0. 

        9 - x² > 0

Factoring,

(3 - x)(3 + x) > 0

We find the critical numbers by setting
each factor = 0

3 - x = 0 gives critical number 3

3 + x = 0 gives critical number -3

Therefore we mark these critical values 
on a number line

-----------o-----------------o----------
 -6 -5 -4 -3 -2 -1  0  1  2  3  4  5  6

Then we test any value of x less than -3, 
to see whether it satisfies the inequality

          (3 - x)(3 + x) > 0

or not.  Let's pick -4 to test

( 3 - (-4) )( 3 + (-4) ) < 0
          (3 + 4)(3 - 4) < 0
                 (7)(-1) < 0
                      -7 < 0

This is false.  So we do not shade the
number line to the left of -3.

-----------o-----------------o----------
 -6 -5 -4 -3 -2 -1  0  1  2  3  4  5  6

Next we test any value of x between -3 and 3,
to see whether it satisfies the inequality

(3 - x)(3 + x) > 0

or not.  Let's pick 0 to test

(3 - 0)(3 + 0) > 0
        (3)(3) > 0
             9 > 0

This is true.  So we DO shade the
number line between -3 and 3.

-----------o=================o----------
 -6 -5 -4 -3 -2 -1  0  1  2  3  4  5  6

Now we test any value of x greater than 3,
to see whether it satisfies the inequality

(3 - x)(3 + x) > 0

or not.  Let's pick 4 to test

(3 - 4)(3 + 4) > 0
       (-1)(7) > 0
            -7 > 0

This is false.  So we do not shade the
number line to the right of 3.

-----------o=================o----------
 -6 -5 -4 -3 -2 -1  0  1  2  3  4  5  6

Now we test the critical numbers themselves
to see whether they satisfy the inequality

(3 - x)(3 + x) > 0

or not.

Testing x = -3, we get 0 > 0 which is true,
so we shade -3

-----------l=================o----------
 -6 -5 -4 -3 -2 -1  0  1  2  3  4  5  6

Testing x = 3, we get 0 > 0 which is true,
so we shade 3

-----------l=================l----------
 -6 -5 -4 -3 -2 -1  0  1  2  3  4  5  6

So the answer is -3 < x < 3

Which in set-builder notation is {x|-3 < x < 3}

or in interval notation [-3, -3]

Edwin